OpenStudy (anonymous):

Kavita wants to borrow an amount of Rs.50000 for 3 years.Bank A charges an compund interest at 8% per annum ,while bank B charges simple interest at 5% semi anually .which offer is more profitable for Kavita.(show the whole solution)

OpenStudy (vishweshshrimali5):

BANK A \[CI = P * (1 + \frac{r}{100})^n\] \[CI = 50000* (1 + \frac{8}{100})^3\] now calculate it BANK B \[SI = \frac{50000*10*3}{100}\] solve both and compare which one is larger that will be ur answer

OpenStudy (anonymous):

but the mode of interest is different

OpenStudy (vishweshshrimali5):

dear, the greater the interest the greater the amount

OpenStudy (anonymous):

i mean to say that Bank A charges it anually and B semi anually

OpenStudy (anonymous):

if bank B charges more will it be more/less profitable

OpenStudy (vishweshshrimali5):

well it will be less profitable if the quantity is more

OpenStudy (anonymous):

hmmm then ur anwer is wrong